How many liters each of a 2% salt solution and an 10% salt solution must be combined to produce 30 liters of a 4% salt solution?
Answer by mangopeeler07(462)

y=liters 2%
x=liters 10%
coefficient=percent

2y+10x=4(30)
2y+10x=120

y+x=30


So you have two equations:
2y+10x=120
y+x=30

Solve for y in the second equation
y=30-x

Plug that into the first equation
2(30-x)+10x=120

Distribute
60-2x+10x=120

Combine like terms
60+8x=120

Subtract 60 from both sides
8x=60

Divide by 8
x=60/8 or 15/2

y=30-x
y=30-15/2
y=60/2-15/2
y=45/2

liters 2%=22.5
liters 10%=7.5